OpenStudy (anonymous):

Find all the roots of the equation (2z-1)^3=1 in the form x+iy.

OpenStudy (anonymous):

I have done it until here. How do I solve it further? \[2\sqrt{2}z ^{3}-6z ^{2}+3\sqrt{2}z-2=0\]

OpenStudy (john_es):

The path you choose to find the solution seems a little hard. I would do something different. Try this, \[(2z-1)^3=1\Rightarrow 2z-1=\theta\] where \theta is the 3rd rooth of 1. So there will be three equations, \[ 2z-1=\theta_1\\ 2z-1=\theta_2\\ 2z-1=\theta_3\] And then solve for z.

OpenStudy (nipunmalhotra93):

do you know the cube roots of unity?

OpenStudy (anonymous):

No, I don't know the cube roots of unity.

OpenStudy (nipunmalhotra93):

by using the polar form of z, you can find all z for which z^3=1....

OpenStudy (science0229):

Have you ever heard of De Moivre's Theorem?

OpenStudy (nipunmalhotra93):

then for those solutions, do what john_es said.

OpenStudy (anonymous):

Yes. I know DeMoivre's Theorem.

OpenStudy (science0229):

It states that \[(\cos \theta+i \sin \theta)^n=\cos n \theta+i \sin n \theta\]

OpenStudy (anonymous):

Okay, but how do I apply it the nth roots of unity?

OpenStudy (science0229):

Do you agree that \[\cos(\theta+2n \pi)=\cos \theta\]and\[\sin(\theta+2n \pi)=\sin \theta\] where n=1,2,3,...

OpenStudy (science0229):

correction: n=0,1,2,3,...

OpenStudy (science0229):

right?

OpenStudy (anonymous):

Yes.

OpenStudy (science0229):

Substitute this into the original De Moivre's Theorem.

OpenStudy (science0229):

Ignore what I just said.

OpenStudy (science0229):

For an equation that looks like x^n=c where x is complex number and c is complex constant, we can apply de moivre's theorem.

OpenStudy (anonymous):

I'm not following you. Can you please do it for one of the equations john posted? :o

OpenStudy (science0229):

sorry. we can write c as following.\[c=a+b i=r(\cos \theta+i \sin \theta)\] where \[r=\sqrt{a^2+b^2}\]and\[\tan \theta=\frac{ b }{ a }\]

OpenStudy (science0229):

From the trig identities...\[c=r(\cos( \theta+2k \pi)+i \sin(\theta+2k \pi))\]

OpenStudy (science0229):

So, \[x^n=r(\cos(\theta+ 2k \pi)+i \sin(\theta+ 2k \pi))\] Take the nth root to get\[x=r^\frac{ 1 }{ n }(\cos\frac{ \theta+2k \pi }{ n }+i \sin\frac{ \theta+2k \pi }{ n })\] where k=0,1,2,...,n-1

OpenStudy (science0229):

There, you have the general solution for x^n=c

OpenStudy (science0229):

The reason why k=0,1,2,..,n-1 is to create 'n' solutions.

OpenStudy (john_es):

May be, @emcrazy14 knows what is the polar form of a complex number. Then, 1 should be put in polar form first. \[1=1_0\] And then do the root, \[\sqrt[n]{a_\theta}=\sqrt[n]{a}_{\theta/n+2\pi k/n}\ \ k=0,1,n-1\] But it is the same as @science0229 said.

OpenStudy (science0229):

Right.

OpenStudy (anonymous):

Okay, so I equate 2z-1 with the first root of unity, then the second and finally the third?

OpenStudy (science0229):

Yeah. Basically.

OpenStudy (science0229):

Also, for future reference, it would be good to memorize the "complex root formula"

OpenStudy (anonymous):

Okay, thank you so much. Actually I didn't do this in class and we were taught the long way of forming the equation, then factorising it and equating to 0 to find the roots.

OpenStudy (science0229):

I forgot: make sure that the theta is between 0 and 2pi

OpenStudy (science0229):

And good luck!